OpenStudy (anonymous):

Give standard and general form to the equation y=0.05x^2 + x + 1

OpenStudy (anonymous):

i am not sure what standard form is, but it might be \(y=a(x-h)^2+k\) in which case your first job it to factor out the \(.05\)

OpenStudy (anonymous):

this might be a little confusing, as the middle term is \(x\) which you have to think of as \(1x\) and since \(1\div .05=100\div 5=20\) you are going to write \[y=.05(x^2+20x)+1\]

OpenStudy (anonymous):

then to write in the form you want, take half of 20 which is 10 and write \[y=.05(x+10)^2+1-\text{something}\] so find the "something" note that \[(x+10)^2=x^2+20x+100\] so \[.05(x+10)^2=.05(x^2+20x+100)=.05x^2+x+5\] but you don't have 5, you have 1 so you have to subtract 4 because \(1-5=-4\) and your final answer is \[y=.05(x+10)^2-4\]

OpenStudy (anonymous):

we can do this another way if you like. we could say the vertex of \[y=..05x^2+x+1\] occurs at \(x=-\frac{b}{2a}=-\frac{1}{2\times .05}=-10\) so we know it looks like \[y=.05(x+10)^2+k\] and to find \(k\) replace \(x\) by 10 in the original expression and get \[k=.05(-10)^2+(-10)+1= .05\times 100-10+1=5-10+1=-4\]

OpenStudy (anonymous):

So y = .05(x + 10)^2 - 4 would be in general form?

OpenStudy (anonymous):

i prefer the second way, because i think it is easiest, but it is up to you yes, i believe that is the "general form" i know it as the vertex form, look in your book to see what "general form" means exactly

OpenStudy (anonymous):

The last part is to find the solution of the function by completing the square

OpenStudy (anonymous):

you mean set it equal to zero and solve right?

OpenStudy (anonymous):

start with \[.05x^2+x+1=0\] and to complete the square the leading coefficient has to be one, so multiply by 20 to get \[x^2+20x+20=0\] subtract 20 to get \[x^2+20x=-20\] then complete the square as before half of 20 is 10, \(10^2=100\) and you write \[(x+10)^2=-20+100=80\]i.e. \[(x+10)^2=80\] take the square root, don't forget the \(\pm\) get \[x+10=\pm\sqrt{80}\] and finally \[x=-10\pm\sqrt{80}\]

OpenStudy (anonymous):

since \(80=16\times 5\) you can also write this as \[x=-10\pm4\sqrt{5}\]

OpenStudy (anonymous):

Thank you I appreciate the help!

OpenStudy (anonymous):

yw